OpenStudy (anonymous):

15. A metal of unknown atomic mass exists as a trivalent ion. When electrolysed in solution using a current of 367mA for 186 min, 0.736g of the metal was deposited on the cathode. Calculate the atomic mass of metal

OpenStudy (anonymous):

@thushananth01

OpenStudy (anonymous):

U have been given the current and time, what can you find from that

OpenStudy (anonymous):

coulombs

OpenStudy (anonymous):

0.367 x 11160 =4095.72 C

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

I coverted the mins into seconds as time is measured in minutes. I converted milliAmps to Amps due to Current been measured in Amps

OpenStudy (anonymous):

yes thats right..Now you can make an equation \[M^{3+}_{(aq)} + 3e^{-} = M _{(s)}\]

OpenStudy (anonymous):

u know that 1 mol of electron is equal to 96500C right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what do i do next??

OpenStudy (anonymous):

Yes 3 mol of electron gives 1 mol of Metal which means 3 * 96500C = 1 Mol of metal so 4095.2C gives how mch mole of metal

OpenStudy (anonymous):

ratio :D

OpenStudy (anonymous):

do you divide 96500C by 4095.2C

OpenStudy (anonymous):

Could you show me how to find the ration please?

OpenStudy (anonymous):

ratio*

OpenStudy (anonymous):



OpenStudy (anonymous):

0.014

OpenStudy (anonymous):

yes u know Mole = mass / Ar u know mole, mass! There you go :)

OpenStudy (anonymous):

so thats the answer = 0.014?

OpenStudy (anonymous):

no ur supposed to find Ar

OpenStudy (anonymous):

use n = mass/Ar where n=no of moles

OpenStudy (anonymous):

52.57 g/mol

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

so thats the answer

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks for your help :)

OpenStudy (anonymous):

no prbs :)